# Tensor Decomposition Networks for Fast Machine Learning Interatomic Potential Computations

### v3

@@ Appendix A Multilinear Maps and Tensor Decomposition in Arbitrary Order
where ${\bm{A}}\in\mathbb{R}^{d_{N+1}\times R}$ and ${\bm{B}}^{(i)}\in\mathbb{R}^{d_{i}\times R}\;(i=1,\dots,N)$. Substituting into Eq.˜4 and rearranging the summation the we obtain
@@ B.1 Error Bound of CP-Based Tensor Product
Theorem B.1 (Equivariance Error Bound of CP Decomposition): Let CG tensor ${\bm{M}}\in\mathbb{R}^{d\times d\times d}$ and $\widehat{{\bm{M}}}$ be the rank-$R$ CP-decomposition-based approximation obtained by Frobenius minimization. For any rotation ${\bm{R}}\in SO(3)$ and any bounded $\rm{SO}(3)$-irreps ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d},\lVert{\bm{x}}\rVert,\lVert{\bm{y}}\rVert\leq C$, we have $\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\;\leq\;2C^{2}\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr)^{1/2},$ where $R_{T}=\lceil R^{1/3}\rceil$ and $\sigma_{k}^{(n)}$ is the $k$-th singular value of mode-$n$ matricization of $\bm{M}$}.
@@
Given any rotation $\bm{R}$} and $\rm{SO}(3)$-irreps ${\bm{x}},{\bm{y}}$,
@@
$\begin{split}\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})&=\lVert\widehat{{\bm{M}}}(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})-D({\bm{R}})\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert\\ &=\lVert\widehat{{\bm{M}}}(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})-{\bm{M}}(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})\\ &\quad+D({\bm{R}}){\bm{M}}({\bm{x}}\otimes{\bm{y}})-D({\bm{R}})\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert\\ &\leq\lVert\widehat{{\bm{M}}}(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})-{\bm{M}}(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})\rVert\\ &\quad+\lVert D({\bm{R}}){\bm{M}}({\bm{x}}\otimes{\bm{y}})-D({\bm{R}})\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert\\ &=\lVert(\widehat{{\bm{M}}}-{\bm{M}})(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})\rVert\\ &\quad+\lVert D({\bm{R}})({\bm{M}}-\widehat{{\bm{M}}})({\bm{x}}\otimes{\bm{y}})\rVert\\ &=\lVert(\widehat{{\bm{M}}}-{\bm{M}})(D({\bm{R}})\otimes D({\bm{R}}))({\bm{x}}\otimes{\bm{y}})\rVert\\ &\quad+\lVert D({\bm{R}})({\bm{M}}-\widehat{{\bm{M}}})({\bm{x}}\otimes{\bm{y}})\rVert\\ &\leq\lVert\widehat{{\bm{M}}}-{\bm{M}}\rVert_{F}\lVert(D({\bm{R}})\otimes D({\bm{R}}))({\bm{x}}\otimes{\bm{y}})\rVert_{F}\\ &\quad+\lVert{\bm{M}}-\widehat{{\bm{M}}}\rVert_{F}\lVert{\bm{x}}\otimes{\bm{y}}\rVert_{F}\\ &=2\lVert\widehat{{\bm{M}}}-{\bm{M}}\rVert_{F}\lVert{\bm{x}}\otimes{\bm{y}}\rVert_{F}\\ &=2\lVert\widehat{{\bm{M}}}-{\bm{M}}\rVert_{F}\lVert{\bm{x}}\rVert\lVert{\bm{y}}\rVert\\ &\leq 2C^{2}\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{2}\Bigr)^{1/2}\end{split}$ (6)
@@ B.2 Universality of CP-Based Tensor Product
Theorem B.2 (Universality of CP Decomposition): Consider $\rm{SO}(3)$-irreps ${\bm{x}}\in V_{1}=\mathbb{R}^{d}$, ${\bm{y}}\in V_{2}=\mathbb{R}^{d_{2}}$ and $V_{3}=\mathbb{R}^{d_{3}}$. For any $\rm{SO}(3)$-equivariant bilinear map $\Phi$, there exist ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$, ${\bm{A}}\in\mathbb{R}^{R\times d_{3}}$ such that $\Phi$ can be written as $\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})\in V_{3},$ if $R$ is sufficiently large.
@@
A bilinear map $\Phi$ is uniquely encoded by a third-order tensor ${\bm{T}}\in V_{3}\otimes V_{1}\otimes V_{2}$ via $\Phi({\bm{x}},{\bm{y}})={\bm{T}}({\bm{x}}\otimes{\bm{y}})$ and all equivariant tensors $\bm{T}$} form the fixed-point subspace $\mathcal{H}=(V_{3}\otimes V_{1}\otimes V_{2})^{\rm{SO}(3)}=\operatorname{Hom}_{\rm{SO}(3)}(V_{1}\otimes V_{2},V_{3})$. Write each irrep $V_{i}$ as a direct sum of irreps
$V_{i}\cong\bigoplus_{\ell=0}^{L}c_{i}^{(\ell)}H^{(\ell)}$
where $c_{i}^{(\ell)}$ is the multiplicity of the irrep of degree $\ell$ in $V_{i}$ and $H^{(\ell)}$ is the $(2\ell+1)$-dimensional irrep. For every admissible path $(\ell_{1},\ell_{2},\ell_{3})$ and multiplicities $1\leq v\leq c_{1}^{(\ell_{1})},1\leq w\leq c_{2}^{(\ell_{2})},1\leq u\leq c_{3}^{(\ell_{3})}$, we pick orthonormal basis vectors ${\bm{a}}_{u}^{(\ell_{3})}\in V_{3}$, ${\bm{b}}_{v}^{(\ell_{1})}\in V_{1}$, ${\bm{c}}_{w}^{(\ell_{2})}\in V_{2}$ and define
${\bm{T}}^{(\ell_{1},\ell_{2},\ell_{3})}_{uvw}={\bm{a}}_{u}^{(\ell_{3})}\otimes{\bm{b}}_{v}^{(\ell_{1})}\otimes{\bm{c}}_{w}^{(\ell_{2})}$
so that $\{{\bm{T}}^{(\ell_{1},\ell_{2},\ell_{3})}_{uvw}\}$ is a basis of $\mathcal{H}$. And then we zero‐pad ${\bm{a}}^{(\ell_{3})}_{u},{\bm{b}}^{(\ell_{1})}_{v},{\bm{c}}^{(\ell_{2})}_{w}$ as the vectors in $V_{3}$, $V_{1}$, and $V_{2}$, respectively. Now let
$\mu=\operatorname{dim}\mathcal{H}=\sum_{\ell_{1}=0}^{L}\sum_{\ell_{2}=0}^{L}\sum_{\ell_{3}=|\ell_{1}-\ell_{2}|}^{\ell_{1}+\ell_{2}\leq L}c_{1}^{(\ell_{1})}c_{2}^{(\ell_{2})}c_{3}^{(\ell_{3})},$
and rename the basis as $\{{\bm{T}}^{(k)}\}_{k=1}^{\mu}$ and vectors ${\bm{a}}^{(\ell_{3})}_{u},{\bm{b}}^{(\ell_{1})}_{v},{\bm{c}}^{(\ell_{2})}_{w}$ as ${\bm{a}}_{k},{\bm{b}}_{k},{\bm{c}}_{k}$ for the corresponding ${\bm{T}}^{(k)}={\bm{T}}^{(\ell_{1},\ell_{2},\ell_{3})}_{uvw}$. Given an arbitrary equivariant tensor $\bm{T}$$\in\mathcal{H}$, we expand it in that basis ${\bm{T}}=\sum_{k=1}^{\mu}\lambda_{k}{\bm{T}}^{(k)}$ and stack the vectors correspondingly as
${\bm{A}}=\bigl[\lambda_{1}{\bm{a}}_{1},\dots,\lambda_{\mu}{\bm{a}}_{\mu}\bigr]^{\!\top},\quad{\bm{B}}=\bigl[{\bm{b}}_{1},\dots,{\bm{b}}_{\mu}\bigr],\quad{\bm{C}}=\bigl[{\bm{c}}_{1},\dots,{\bm{c}}_{\mu}\bigr].$
For all ${\bm{x}},{\bm{y}}$ one then has $\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\!\top}{\bm{x}}\odot{\bm{C}}^{\!\top}{\bm{y}}).$ Taking $R\geq\mu$ supplies enough columns/rows to hold this equation; any surplus columns can be zero-padded. ∎
@@ Appendix D Additional Ablation Studies
Ablation study of CP decomposition rank.
To further demonstrate the practical implication of CP-decomposition-based tensor product and the adopted scheduler, we conduct the ablation study of ranks over $n$-body system dataset [41]. In Table˜7 of the n-body system experiment, the TDN is trained with a maximum angular degree $L=2$, three interaction blocks, and a hidden dimension of 72 in 5000 epochs over a single NVIDIA RTX 2080Ti-11GB GPU. As shown in the table, our adopted schedule uses much smaller ranks yet matches the accuracy obtained with the largest ranks, corresponding to the results of rank schedule selection in. Note that the largest rank is deduced from the upper bound in Section˜3.3.
$R$ $n$-Body Testing MSE Equivariance Error
10 0.0081 0.60
15 0.0064 0.42
20 0.0051 0.21
28 (Our scheduler) 0.0040 0.02
81 (Highest rank) 0.0039 < 0.01
TDN w/o CP + PS 0.0038 -
@@
Table˜8 presents the time ablation study of path-weight sharing and CP decomposition over the TDN model. As shown in the table, CP decomposition significantly increases the GPU and CPU throughput while path-weight sharing mechanism substantially reduces the parameter count of the model. Note that because TDN removes the depth-wise tensor-product operator, a TDN without CP decomposition and without path-weight sharing in the tensor-product and equivariant-linear layers is not identical to the vanilla Equiformer. All experiments are run on a single NVIDIA A100-80GB GPU and Xeon Gold 6258R processor with a batch size of 128. All experiments are conducted under identical irrep configurations across varying $L$ with 256 irrep-channel dimension, six graph transformer layers, and 8 attention heads of each layer.

### v4

@@ Appendix A Multilinear Maps and Tensor Decomposition in Arbitrary Order
where ${\bm{A}}\in\mathbb{R}^{d_{N+1}\times R}$ and ${\bm{B}}^{(i)}\in\mathbb{R}^{d_{i}\times R}\;(i=1,\dots,N)$. Substituting into Eq.˜5 and rearranging the summation the we obtain
@@ B.1 Error Bound of CP-Based Tensor Product
Theorem B.1 (Equivariance Error Bound of CP Decomposition): Let CG tensor ${\bm{M}}\in\mathbb{R}^{d\times d\times d}$ and $\widehat{{\bm{M}}}$ be the rank-$R$ CP-decomposition-based approximation obtained by Frobenius minimization. For any rotation ${\bm{R}}\in SO(3)$ and any bounded representations ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d},\lVert{\bm{x}}\rVert,\lVert{\bm{y}}\rVert\leq C$, we have $\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\;\leq\;2C^{2}\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr)^{1/2},$ where $R_{T}=\lfloor R^{1/3}\rfloor$ and $\sigma_{k}^{(n)}$ is the $k$-th singular value of mode-$n$ matricization of $\bm{M}$}.
@@
Given any rotation $\bm{R}$} and $\rm{SO}(3)$-representations ${\bm{x}},{\bm{y}}$,
@@
[…]
Let ${\bm{M}}_{\text{truncated}}$ be the truncated Tucker approximation of $\bm{M}$} with multilinear ranks $(R_{T},R_{T},R_{T})$. A priori approximation error bound [32] gives
$\bigl\lVert{\bm{M}}-{\bm{M}}_{\text{truncated}}\bigr\rVert_{F}\;\leq\;\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr)^{1/2}.$
CP-Decomposing the Tucker core with a size of $(R_{T},R_{T},R_{T})$ yields CP rank at most $R_{T}^{3}$. Therefore, the truncated Tucker tensor ${\bm{M}}_{\text{truncated}}$ can be written as a CP tensor $\widehat{{\bm{M}}}$ with rank at most $R_{T}^{3}$; in other words, there exists a CP tensor $\widehat{{\bm{M}}}$ such that $\text{rank}_{\text{CP}}(\widehat{{\bm{M}}})\leq R_{T}^{3}$ and
$\bigl\lVert{\bm{M}}-\widehat{{\bm{M}}}\bigr\rVert_{F}\;\leq\;\bigl\lVert{\bm{M}}-{\bm{M}}_{\text{truncated}}\bigr\rVert_{F}\;\leq\;\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr)^{1/2}.$
Inserting this to Eq.˜7 yields
$\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\leq 2C^{2}\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr)^{1/2}.$
@@ B.2 Universality of CP-Based Tensor Product
Theorem B.2 (Universality of CP Decomposition): Consider $\rm{SO}(3)$-representations ${\bm{x}}\in V_{1}\cong\mathbb{R}^{d_{1}}$, ${\bm{y}}\in V_{2}\cong\mathbb{R}^{d_{2}}$ and co-domain $V_{3}\cong\mathbb{R}^{d_{3}}$. For any $\rm{SO}(3)$-equivariant bilinear map $\Phi$, there exist ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$, ${\bm{A}}\in\mathbb{R}^{d_{3}\times R}$ such that $\Phi$ can be written as $\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})\in V_{3},$ with $R\leq d_{1}d_{2}$.
@@
A bilinear map $\Phi$ is uniquely encoded by a third-order tensor ${\bm{T}}\in V_{3}\otimes V_{1}^{*}\otimes V_{2}^{*}$ via $\Phi({\bm{x}},{\bm{y}})={\bm{T}}({\bm{x}}\otimes{\bm{y}})$ and all equivariant tensors $\bm{T}$} form the subspace $\mathcal{H}=(V_{3}\otimes V_{1}^{*}\otimes V_{2}^{*})^{\rm{SO}(3)}=\operatorname{Hom}_{\rm{SO}(3)}(V_{1}\otimes V_{2},V_{3})$. For simplicity, we consider the multiplicity-free case $V_{i}\cong\bigoplus_{\ell=0}^{L}H^{(\ell)}$, and the general case replaces the scalars below by linear maps on multiplicity spaces. Decomposing $\mathcal{H}$ into irreps gives
$\mathcal{H}\;\cong\;\bigoplus_{\ell_{1},\ell_{2},\ell_{3}}\mathrm{Hom}_{\mathrm{SO}(3)}\!\bigl(H^{(\ell_{1})}\otimes H^{(\ell_{2})},\,H^{(\ell_{3})}\bigr).$
By the Clebsch-Gordan decomposition, $\mathrm{Hom}_{\mathrm{SO}(3)}(H^{(\ell_{1})}\otimes H^{(\ell_{2})},H^{(\ell_{3})})$ is one-dimensional when $|\ell_{1}-\ell_{2}|\leq\ell_{3}\leq\ell_{1}+\ell_{2}$ and zero otherwise. For each path $(\ell_{1},\ell_{2},\ell_{3})$, there is a unique map $C_{\ell_{1},\ell_{2}}^{\ell_{3}}:H^{(\ell_{1})}\times H^{(\ell_{2})}\to H^{(\ell_{3})}$ with matrix elements the Clebsch-Gordan coefficients $C^{\ell_{3},m_{3}}_{\ell_{1},m_{1},\ell_{2},m_{2}}$, as presented by Eq.˜1. Therefore, any equivariant bilinear map $\Phi$ admits the expansion
$\Phi\;=\;\sum_{\ell_{1},\ell_{2},\ell_{3}}\alpha_{\ell_{1},\ell_{2}}^{\ell_{3}}\,C_{\ell_{1},\ell_{2}}^{\ell_{3}},$
for some scalars $\alpha_{\ell_{1},\ell_{2}}^{\ell_{3}}$. Now we index coordinates of $V_{1},V_{2},V_{3}$ by $i=(\ell_{1},m_{1})$, $j=(\ell_{2},m_{2})$, and $k=(\ell_{3},m_{3})$. Let the third-order tensor ${\bm{T}}_{kij}\;=\;\alpha_{\ell_{1},\ell_{2}}^{\ell_{3}}\,C^{\ell_{3},m_{3}}_{\ell_{1},m_{1},\ell_{2},m_{2}}$ and then $\Phi({\bm{x}},{\bm{y}})_{k}=\sum_{i,j}{\bm{T}}_{kij}\,{\bm{x}}_{i}\,{\bm{y}}_{j}$. To write $\Phi$ in the CP decomposition form, we take $R=d_{1}d_{2}$ and index the column dimension by pairs $(i,j)$. Define ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$, and ${\bm{A}}\in\mathbb{R}^{d_{3}\times R}$ by
${\bm{B}}_{i^{\prime},(i,j)}=\mathbf{1}[i^{\prime}=i],\qquad{\bm{C}}_{j^{\prime},(i,j)}=\mathbf{1}[j^{\prime}=j],\qquad{\bm{A}}_{k,(i,j)}={\bm{T}}_{kij}.$
with $\mathbf{1}[\cdot]$ the indicator function. Then $({\bm{B}}^{\top}{\bm{x}})_{(i,j)}={\bm{x}}_{i}$ and $({\bm{C}}^{\top}{\bm{y}})_{(i,j)}={\bm{y}}_{j}$, $({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})_{(i,j)}={\bm{x}}_{i}{\bm{y}}_{j}$, and therefore
$\bigl({\bm{A}}({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})\bigr)_{k}=\sum_{i,j}{\bm{T}}_{kij}\,{\bm{x}}_{i}\,{\bm{y}}_{j}=\Phi({\bm{x}},{\bm{y}})_{k}.$
For larger $R$, one can pad ${\bm{A}},{\bm{B}},{\bm{C}}$ with zero columns. In practice, we approximate $\bm{T}$} by a lower-rank CP decomposition with computed ${\bm{A}},{\bm{B}},{\bm{C}}$ $R\ll d_{1}d_{2}\propto(L+1)^{4}$, corresponding to Fig.˜2. ∎
@@ Appendix D Additional Ablation Studies
Table˜7 presents the time ablation study of path-weight sharing and CP decomposition over the TDN model. As shown in the table, CP decomposition significantly increases the GPU and CPU throughput while path-weight sharing mechanism substantially reduces the parameter count of the model. Note that because TDN removes the depth-wise tensor-product operator, a TDN without CP decomposition and without path-weight sharing in the tensor-product and equivariant-linear layers is not identical to the vanilla Equiformer. All experiments are run on a single NVIDIA A100-80GB GPU and Xeon Gold 6258R processor with a batch size of 128. All experiments are conducted under identical irrep configurations across varying $L$ with 256 irrep-channel dimension, six graph transformer layers, and 8 attention heads of each layer.
@@
Ablation study of CP decomposition rank.
To further demonstrate the practical implication of CP-decomposition-based tensor product and the adopted scheduler, we conduct the ablation study of ranks over $n$-body system dataset [44]. In Table˜8 of the n-body system experiment, the TDN is trained with a maximum angular degree $L=2$, three interaction blocks, and a hidden dimension of 72 in 5000 epochs over a single NVIDIA RTX 2080Ti-11GB GPU. As shown in the table, our adopted schedule uses much smaller ranks yet matches the accuracy obtained with the largest ranks, corresponding to the results of rank schedule selection in. Note that the largest rank is deduced from the upper bound in Section˜3.3.
$R$ $n$-Body Testing MSE Equivariance Error
10 0.0081 0.60
15 0.0064 0.42
20 0.0051 0.21
28 (Our scheduler) 0.0040 0.02
81 (Highest rank) 0.0039 < 0.01
TDN w/o CP + PS 0.0038 -
Ablation of CP-Decomposition-Based Tensor Product for TDN and NequIP.
Under the settings described in Section˜4.6, Table˜9 shows that for TDN trained for 60 epochs, enabling both mechanisms yields slightly higher validation energy MAE and force RMSE while reducing training time by roughly 78%, suggesting only a minor effect on predictive accuracy. We also evaluate NequIP [9] trained for 100 epochs. Adding the path-weight sharing tensor product and CP decomposition preserves downstream accuracy while cutting computational cost nearly 74%. Note that accuracy metrics are reported from abbreviated training runs intended only to assess relative trends; the fully tuned and converged results are reported in Table˜1.
Table 9: Ablation study of the path-weight sharing tensor product (PS) and CP-decomposition-based tensor product (CP). Results are shown for TDN and NequIP [9] on PubChemQCR-S dataset.
Model Validation
Energy MAE Force RMSE Training Time
(meV/atom) $\downarrow$ (meV/Å) $\downarrow$ (min/epoch) $\downarrow$
TDN w/o CP + PS 12.74 92.74 19.0
TDN 12.95 96.53 4.2 ($\times$ 0.22)
NequIP 11.22 90.09 7.5
NequIP + CP + PS 11.15 92.49 2.0 ($\times$ 0.26)